# Multilevel Fast Multipole Algorithm for Electromagnetic Scattering of Metasurfaces using a Static Mode Representation

### v1

@@ V.2 Computational costs
Figure 4: (a) CPU time required for the PMCHWT system solution, comparing the RWG with Static Mode Expansion. The code is implemented in FORTRAN and runs on a single CPU. (b) Memory requirement comparison between the RWG with Static Mode Expansion. The $\mathcal{O}(p\log p)$ (continuous line) and $\mathcal{O}\left(p^{1.25}\log p\right)$ (dashed line) curves are also plotted for reference. (Inset) Surface mesh used for the computational complexity test.
@@
We conducted a complexity study on metasurfaces consisting of a varying number up to $5\cdot 10^{4}$ of identical gold nano-spheres, each of radius R = 100 nm. The particles are arranged according to a golden-angle spiral pattern. This aperiodic pattern is particularly convenient for our study as it is deterministic, and allows us to increase the dimension of the array one particle at a time without appreciably changing the array’s density. The array is excited by a linearly polarized plane wave of wavelength $\lambda=600$ nm, propagating along the axis orthogonal to the array’s plane. Each sphere has been discretized by a surface mesh, shown in the inset of Fig. 4 (a), with $100$ nodes, $196$ triangles, $294$ edges, which results in an average edge length approximately equal to $\lambda/15$. In Fig 4, we present a comparison between the CPU time and memory requirements using the MLFMA with RWG basis function and static mode expansion with $N^{\|}=N^{\perp}=10$. Simulation results show that the use of MLFMA, in both cases RWG and SME, reduces the computational complexity of the matrix-vector product from $\mathcal{O}(p^{2})$ to $\mathcal{O}(p\log p)$, as expected. The number of degrees of freedom, and the computational time of each case study reported in Fig. 4 are listed in Tab. 1. In particular, with the use of the SME, this simulation is approximately $3\div 5$ times faster.
@@
However, the speed-up depends on various factors, such as the number of points used for interpolation/anterpolation between different levels of the quadtree, or the number of points used to sample the transfer function $T_{L}$ on the unit sphere. Above all, mesh density on each particle plays a very important role. Indeed, the advantages introduced by the MLFMA-SME become much more apparent with denser meshes, necessary for discretizing irregular objects, such as nanofins and nanorods typically used in metalens. The reason why the speed-up is less than the number of unknowns’ compression ratio is attributable to the $k$-space representation of the MLFMA operators, which means that the computational complexity is not only linked to the number of unknowns but also to the electrical size of the array.
Table 1: Comparison between RWG and SME.
@@
Another crucial advantage introduced by SME in the MLFMA concerns memory requirements. Although the translation function $T_{L}$ is the same in both RWG and SME, as it doesn’t depend on the basis function, the matrices representing the radiation and receive operators are $\kappa$ times smaller, and the self impedance matrix $Z_{s}$ is $\kappa^{2}$ times smaller, where $\kappa=n_{e}/(N^{\parallel}+N^{\perp})$ is the total number of unknowns’ compression ratio. Additionally, after $m$ iterations, computing the GMRES solution requires storing all the basis vectors of the Krylov subspace $\mathcal{K}_{m}$, leading to a memory requirement of $\mathcal{O}(mN)$, where $N$ is the total number of unknowns. As $N$ increases, the storage cost may become prohibitive, even for a small number of iterations $m$. As previously mentioned, SME can drastically reduce $N$, allowing for the simulation of larger structures that were impracticable before.
Figure 5: (a) Electric field magnitude on the particles’ surface of a finite Vogel spiral composed of $40\cdot 10^{3}$ gold bricks of dimension $200\times 200\times 100$ nm under the excitation of a linearly polarized plane wave at wavelength $\lambda$ = 600 nm. The array size is about $70\mu$m, which is roughly $115\lambda$. The solution has been computed by using the MLFMA-SME with $N^{\|}=N^{\perp}=10$. (b) Electric field magnitude on the particles’ surface of a Moiré superlattice array composed of 20160 gold nanodisks of radius $100$ nm under the excitation of a linearly polarized plane wave at wavelength $\lambda$ = 600 nm. The array size is about $46\times 53\mu$m, which is roughly $(76\times 87)\lambda$. The solution has been computed by using the MLFMA-SME with $N^{\|}=N^{\perp}=10$. (c) Electric field magnitude on the particles’ surface of a canonical metalens, lying on the xy plane, and composed of $10^{4}$ $\mathrm{TiO}_{2}$ nanofins of dimension $410\times 85\times 600$ nm under the excitation of a circularly polarized plane wave at wavelength $\lambda=660$ nm. The array size is $66\,\mu$m, which is $100\lambda$. The solution has been computed by using the MLFMA-SME with $N^{\|}=N^{\perp}=10$. (d) Electric field normalized intensity on the $x-z$ cross-sectional plane. The achieved focal length is consistent with the prescribed value of $f=37.5\lambda$.
V.3 Electrically large particle arrays
The use of the MLFMA-SME finds its natural application in the numerical solution of the electromagnetic scattering problem by large metasurfaces, where an array, whose dimension can be much larger than the incident wavelength, is made of $p$ objects with identical shape but with different orientation and size. We show three exemplificative case studies of the application of this method to metasurfaces and metalenses: a golden angle spiral of gold bricks, and a Moirè lattice of gold disks and a canonical metalens of $\text{TiO}_{2}$ nano-fins.
@@ V.3 Electrically large particle arrays
We first consider a finite-size golden angle spiral array having linear dimension of $115\lambda$. The array is made of $40\cdot 10^{3}$ gold bricks of dimension $200\times 200\times 100$ nm. We model this shape as a superellipsoid, using Eq. (38) with $m=n=0.25$, $a_{x}=a/2$, $a_{y}=a_{z}=a$. Each particle is described by a surface mesh having $436$ nodes, $868$ triangles, and $1302$ edges. The $i$-th particle is rotated with respect to $\hat{\mathbf{x}}$ by an angle $\theta_{i}=2\pi(i-1)/p$. The array is excited by a plane wave, linearly polarized in the plane of the array, along the $\hat{\mathbf{x}}$-axis, and propagating in the direction orthogonal to the array’s plane with wavelength $\lambda=600$ nm. The array size is about $70\mu$m. In Fig. 5a, we show the electric field magnitude on the particles’ surface. The solution is computed by using the MLFMA-SME as detailed in this work using the static mode expansion with $N^{\|}=N^{\perp}=10$, i.e. 40 unknowns per particle. The total computational time is about 20.4 hours on a single CPU.
@@
Next, we now consider an example of 2.5D metastructures, which are stacked layers of interacting metasurface layers. They provide sufficient degrees of freedom to implement efficient multifunctional devices. Specifically, we consider a two-layer Moiré superlattice array of $20160$ gold nanodisks of radius $R=100$ nm. We model this shape as a superellipsoid, whose boundary has the implicit equation (38), with $m=1$, $n=0.25$, $a_{x}=a_{y}=R$, $a_{z}=R/2$. Each particle is described by a surface mesh having $588$ nodes, $1172$ triangles, and $1758$ edges. The centers of the particles are arranged at the vertices of two layers, obtained by periodically replicating a hexagon with a side length of $400$ nm. The second layer is located at a height $h=300$ nm above the first and is rotated by an angle $\theta=5$° with respect to it. The array is excited by a plane wave, linearly polarized in the plane of the array, along the $\hat{\mathbf{x}}$-axis, and propagating in the direction orthogonal to the array’s plane with wavelength $\lambda=600$ nm. The array size is about $46\times 53\,\mu$m, which is roughly $(76\times 87)\lambda$. In Fig. 5b we show the electric field magnitude on the particles’ surface, while in Fig. 2 we show the longitudinal and transverse modes. The solution is computed using the MLFMA-SME with $N^{\|}=N^{\perp}=10$. The total computational time is less than 6 hours on a single CPU.
@@
Eventually, we consider an example of a canonical metalens [13], composed of $10^{4}$ $\mathrm{TiO}_{2}$ nanofins of dimension $410\times 85\times 600$ nm. We model this shape as a superellipsoid, whose boundary has the implicit equation (38), with $m=n=0.25$. The total computational time is about 15 hours on a single CPU. Each particle is described by a surface mesh having $198$ nodes, $392$ triangles, and $588$ edges. To achieve the desired phase profile, the $i$-th particle, centered at $(x_{i},y_{i})$, is rotated with respect to $\hat{\mathbf{x}}$ by an angle [13, 44]
@@
where $f$ is the focal length. The metalens has a desired numerical aperture (NA) of 0.8, resulting in a focal length $f=37.5\lambda$. The array is excited by a plane wave, circularly polarized in the plane of the array and propagating in the direction orthogonal to the array’s plane with wavelength $\lambda=660$ nm. The array size is $66\times 66\,\mu$m, which is $(100\times 100)\lambda$. In Fig. 5c we show the electric field magnitude on the particles’ surface. The solution is computed using the MLFMA-SME with $N^{\|}=N^{\perp}=10$. In Fig. 5d we show the normalized intensity $\|\mathbf{E}(x,z)\|^{2}$ on the $x-y$ cross-sectional plane. The achieved focal length is consistent with the prescribed value.
@@ VI Conclusion
In this work, we proposed a fast integral equation solver suitable for large particle arrays, even those hundreds of wavelengths in diameter, comprising numerous arbitrarily shaped objects with varying orientations and sizes. These scenarios are commonly encountered in the electromagnetic modeling of metasurfaces and metalenses. Our approach combines the Multilevel Fast Multipole Algorithm (MLFMA) with the PMCHWT formulation, utilizing a particular class of entire domain basis function defined on the isolated particle, termed Static Mode [18].
To validate the accuracy of the proposed method, we analyze the scattering from finite aperiodic arrays of nano-spheres and compare the results with those obtained using the multiparticle Mie theory [19]. We report the CPU-time and memory requirements as functions of the number of particles, comparing the static mode expansion to a more traditional approach using RWG basis functions. Our findings indicate that the MLFMA-SME is $3$ to $5$ times faster and requires about $5$ times less memory than the traditional MLFMA employing RWG basis functions. Furthermore, this speed and memory advantages significantly improve as the density of the surface mesh increases.
Additionally, we demonstrate the capabilities of the MLFMA-SME method on different arrays: a golden-angle spiral array with a diameter of $115\lambda$, consisting of 40,000 gold bricks; a two-layer Moiré superlattice made up of 20,160 gold nanodisks, with overall dimensions of $76\lambda\times 87\lambda$; a canonical metalens composed of 10,000 $\mathrm{TiO}_{2}$ nanofins, covering an area of $100\lambda\times 100\lambda$ and having focal length $f=37.5\lambda$.
Our results show that SME drastically reduces the total number of unknowns, leading to significant savings in CPU time and memory requirements for the numerical solution of scattering problems from arrays of identical particles compared to the classical MLFMA implementation with sub-domain basis functions. This advancement enables the simulation of large structures that were previously impracticable.

### v2

@@ V.2 Computational time and memory cost
Figure 4: (a) CPU time required for the MLFMA solution of the PMCHWT system using RWG with static mode representation. The code runs on a single CPU. (b) Memory requirement comparison between the RWG with static mode representation. The $\mathcal{O}(p\log p)$ (continuous line) and $\mathcal{O}\left(p^{1.25}\log p\right)$ (dashed line) curves are also plotted for reference. (Inset) Surface mesh used for the computational complexity test.
@@
We conducted a complexity study on metasurfaces consisting of a varying number up to $5\cdot 10^{4}$ of identical gold nano-spheres, each of radius R = 100 nm. The particles are arranged according to the golden-angle spiral pattern, defined in Sec. V.1. The array is excited by a linearly polarized plane wave of wavelength $\lambda=600$ nm, propagating along the direction orthogonal to the array’s plane. Each sphere has been discretized by the surface mesh shown in the inset of Fig. 4 (a), with $100$ nodes, $196$ triangles, $294$ edges, which results in an average edge length approximately equal to $\lambda/15$.
In Fig. 4, we compare the CPU time and memory requirements for simulations using the multilevel fast multipole algorithm with RWG basis functions, denoted as MLFMA-RWG, and with the static modes basis, denoted as MLFMA-SMR, with $N^{|}=N^{\perp}=10$. The results demonstrate that, in both cases of RWG and SMR, the MLFMA reduces the computational complexity of the matrix-vector product from $\mathcal{O}(p^{2})$ to $\mathcal{O}(p\log p)$. The degrees of freedom and the computational time for each case are detailed in Tab. 1. Notably, the SMR-based computation is approximately 3 to 5 times faster than the corresponding computation using RWG basis functions. It is important to highlight that, for both RWG and SMR, the error compared to Mie theory is comparable, with an approximate error $\varepsilon\left\{C_{sca}\right\}\approx 3\%$.
@@
However, the speed-up depends on various factors, including the number of points used for interpolation/anterpolation between different levels of the quadtree, or the number of points used to sample the transfer function $T_{L}$ on the unit sphere. Above all, mesh density on each particle plays a very important role. Indeed, the advantages introduced by the MLFMA-SMR become much more apparent with denser meshes, necessary for discretizing irregular objects, such as nanofins and nanorods typically used in metalens.
In the investigated example, the speed-up is less than the number of unknowns’ compression ratio $\kappa=n_{e}/(N^{\parallel}+N^{\perp})$ between the number of basis function used in the RWG and the SMR cases. This fact is attributable to the computational time required by $k$-space representation of the MLFMA operators, which means that the computational complexity is not only linked to the number of unknowns but also to the electrical size of the array.
Table 1: Comparison between RWG and SMR.
@@
Figure 5: (a) Surface mesh of a brick, consisting of 436 nodes, 868 triangles, and 1302 edges. (b) Surface mesh of a nanodisk, consisting of 588 nodes, 1172 triangles, and 1758 edges. (c) Surface mesh of a nanofins, consisting of 468 nodes, 932 triangles, and 1398 edges.
Another key advantage of using the SMR in the MLFMA is the significant reduction in memory requirements. Although the translation function $T_{L}$ is the same in both RWG and SMR, as it doesn’t depend on the basis function, the matrices representing the radiation and receiving patterns are reduced in size by a factor of $\kappa$, the self-impedance matrix $Z_{s}$ is reduced by $\kappa^{2}$, where $\kappa$ is the compression ratio.
Figure 6: (a) Electric field magnitude on the particles’ surface of a finite Vogel spiral composed of $40\cdot 10^{3}$ gold bricks of dimension $200\times 200\times 100$ nm under the excitation of a linearly polarized plane wave at wavelength $\lambda$ = 600 nm. The array size is roughly $115\lambda$. (b) Electric field magnitude on the particles’ surface of a Moiré superlattice array composed of 20160 gold nanodisks of radius $100$ nm under the excitation of a linearly polarized plane wave at wavelength $\lambda$ = 600 nm. The array size is roughly $(76\times 87)\lambda$. (c) Electric field magnitude on the particles’ surface of a canonical metalens, lying on the xy plane, and composed of $10^{4}$ $\mathrm{TiO}_{2}$ nanofins of dimension $410\times 85\times 600$ nm under the excitation of a circularly polarized plane wave at wavelength $\lambda=660$ nm. The array size is $100\lambda$. In all three cases investigated, the solution has been computed by using the MLFMA-SMR with $N^{\|}=N^{\perp}=10$.
Additionally, the memory needed to compute the GMRES solution after $m$ iterations is proportional to the size of the Krylov subspace $\mathcal{K}_{m}$, leading to a memory requirement of $\mathcal{O}(mN)$, where $N$ is the total number of unknowns. As $N$ increases, the storage cost can become prohibitive, even for a relatively small number of iterations $m$. The SMR method significantly reduces $N$, enabling the simulation of larger structures that were previously infeasible due to memory constraints.
V.3 Electrically large metasurfaces
The use of the MLFMA-SMR finds its natural application in the numerical solution of the electromagnetic scattering problem by large metasurfaces, where an array, whose dimension can be much larger than the incident wavelength, is made of $p$ objects with identical shape but with different orientation and size. We show three exemplificative case studies: a golden angle spiral of gold bricks, and a Moirè lattice of gold disks, and a canonical metalens of $\text{TiO}_{2}$ nano-fins.
@@ V.3 Electrically large metasurfaces
We first consider a finite-size golden angle spiral array, as defined in Sec. V.1, having linear dimension of $115\lambda$. The array is made of $40\cdot 10^{3}$ gold bricks of dimension $200\times 200\times 100$ nm. We model this shape as a superellipsoid, using Eq. (32) with $m=n=0.25$, $a_{x}=a/2$, $a_{y}=a_{z}=a$. Each particle is described by the surface mesh shown in Fig. 5 (a). The $i$-th particle is rotated with respect to $\hat{\mathbf{x}}$ by an angle $\theta_{i}=2\pi(i-1)/p$. The array is excited by a plane wave, linearly polarized in the plane of the array, along the $\hat{\mathbf{x}}$-axis, and propagating in the direction orthogonal to the array’s plane with wavelength $\lambda=600$ nm. The array size is about $70\mu$m. In Fig. 6(a), we show the electric field magnitude on the particles’ surface. The solution is computed by using the MLFMA-SMR using the static mode representation with $N^{\|}=N^{\perp}=10$, which corresponds to 40 unknowns per particle. The total computational time is about 20.4 hours on a single CPU. The GMRES required approximately 15 hours and converged after 921 iterations.
@@
Next, we now consider an example of 2.5D metastructures, which are stacked layers of interacting objects. They provide sufficient degrees of freedom to implement efficient multifunctional devices. Specifically, we consider a two-layer Moiré superlattice made by a total number of $20160$ gold nanodisks of radius $R=100$ nm and height $50$ nm. We model the nanodisk as a superellipsoid, whose boundary has the implicit equation (32), with $m=1$, $n=0.25$, $a_{x}=a_{y}=R$, $a_{z}=R/2$. Each particle is described by the surface mesh shown in Fig. 5 (b). The centers of the particles are arranged at the vertices of two layers, obtained by periodically replicating a hexagon with a side length of $400$ nm. The second layer is located at a height $h=300$ nm above the first and is rotated by an angle $\theta=5^{\circ}$ with respect to it. The array is excited by a plane wave, linearly polarized in the plane of the array, along the $\hat{\mathbf{x}}$-axis, and propagating in the direction orthogonal to the array’s plane with wavelength $\lambda=600$ nm. The array size is about $46\times 53\,\mu$m, which is roughly $(76\times 87)\lambda$. The solution is computed using the MLFMA-SMR with $N^{\|}=N^{\perp}=10$. The total computational time is less than 6 hours on a single CPU. The GMRES required approximately 5.2 hours and converged after 564 iterations.
@@
Eventually, we consider an example of a canonical metalens [29], composed of $10^{4}$ $\mathrm{TiO}_{2}$ nanofins of dimension $410\times 85\times 600$ nm. We model this shape as a superellipsoid, whose boundary has the implicit equation (32), with $m=n=0.25$. Each particle is described by the surface mesh shown in Fig. 5 (c). To achieve the desired phase profile, the $i$-th particle, centered at $(x_{i},y_{i})$, is rotated with respect to $\hat{\mathbf{x}}$ by an angle [29, 68]
@@
where $f$ is the focal length. The metalens has a desired numerical aperture (NA) of 0.8, resulting in a focal length $f=37.5\lambda$. The array is excited by a plane wave, circularly polarized in the plane of the array and propagating in the direction orthogonal to the array’s plane with wavelength $\lambda=660$ nm. The array size is $66\times 66\,\mu$m, which is $(100\times 100)\lambda$. The solution is computed using the MLFMA-SMR with $N^{\|}=N^{\perp}=10$. The total computational time is about 16.8 hours on a single CPU. The GMRES required approximately 10.6 hours and converged after 1019 iterations. In Fig. 6(c) we show the electric field magnitude on the particles’ surface. In Fig. 7(d) we show the normalized intensity $\|\mathbf{E}(x,z)\|^{2}$ on the $x$-$y$ cross-sectional plane. The achieved focal length is consistent with the prescribed value.
Figure 7: Electric field normalized intensity on the $x-z$ cross-sectional plane. The achieved focal length is consistent with the prescribed value of $f=37.5\lambda$.
@@ VI Conclusion
In this work, we proposed a fast integral equation solver suitable for large particle arrays, even hundreds of wavelengths in diameter, comprising arbitrarily shaped objects with varying orientations and sizes. These scenarios are commonly encountered in the electromagnetic modeling of metasurfaces and metalenses. Our approach combines the Multilevel Fast Multipole Algorithm (MLFMA) with the Poggio-Miller-Chang-Harrington-Wu-Tsai (PMCHWT) formulation, utilizing a particular class of entire domain basis functions defined on the isolated particle, termed static modes [35]. The static mode representation (SMR) significantly reduces the number of unknowns compared to discretizations based on sub-domain basis functions, such as RWG, thereby lowering both CPU time and memory requirements.
We validated the accuracy of the proposed MLFMA-SMR method by analyzing the scattering from finite aperiodic arrays of nano-spheres and comparing the results with those obtained using the multiparticle Mie theory [36]. Then, we report the CPU-time and memory requirements as functions of the number of particles composing the array, comparing the static mode representation to a more traditional approach using RWG basis functions. Our findings indicate that the MLFMA-SMR is $3$ to $5$ times faster and requires about $5$ times less memory than the traditional MLFMA employing RWG basis functions, while maintaining comparable accuracy. Furthermore, these speed and memory advantages significantly improve as the density of the surface mesh increases.
Eventually, we showcase the capabilities of the MLFMA-SMR method on different arrays: a golden-angle spiral array with a diameter of $115\lambda$, consisting of 40,000 gold bricks; a two-layer Moiré superlattice made up of 20,160 gold nanodisks, with overall dimensions of $76\lambda\times 87\lambda$; and a canonical metalens composed of 10,000 $\mathrm{TiO}_{2}$ nanofins, covering an area of $100\lambda\times 100\lambda$ and having a focal length $f=37.5\lambda$.
Our results demonstrate that the MLFMA-SMR significantly reduces the total number of unknowns, leading to substantial savings in CPU time and memory requirements for the numerical solution of scattering problems from arrays of identical particles, compared to the classical MLFMA implementation with sub-domain basis functions. This advancement makes it feasible to simulate and design large metasurfaces.